sqrt(x-6)+3=sqrt(x)
does this have a real solution?
first i subtracted the 3 over, squared both sides and ended up with x=(25/4) but when i plug that in, it was not a true statement.
then i started over and squared both sides like the original problem shows and got the square root equal to a negative so that would be "no real solution".
thank you for looking at this!!

Found 2 solutions by jim_thompson5910, scott8148:
Answer by jim_thompson5910(35256)

You are correct. No real solutions exist.

Answer by scott8148(6628)

square roots can be positive or negative

25/4 - 6 = 1/4 ___ sqrt(1/4) = ± 1/2

-1/2 + 3 = 5/2

sqrt(25/4) = ± 5/2